A bullet loses 10% of its speed after passing through a wooden plank.How  many  such planks should be placed one after the another so that bullet does not come out?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear neeru

let the thickness of one block is t

(9v/10)2 = v2 - 2at

.81v2 = v2 - 2at

2at = .19v2.........................1

let n block is used

0 = v2 - 2a(nt)

n = v2 /2at

  = v2 /.19v2

  = 100/19

  = 5.26


so 6 block is used

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin